OpenStudy (anonymous):

The volume of a rectangular box is equal to the length times the width times the height. Write an expression for the volume of a box with length 15 cm, width w cm, and height 10 cm. Be sure to simplify the expression. A. 45 + 3w B. 150w C. 30 + 2w D. 150w2

OpenStudy (anonymous):

V = l*w*h l = 15 w = w h = 10 V = 15*w*10 = 150w

OpenStudy (anonymous):

thanks for the help. i needed it